metallic spherical shell of radius 46.6 cm has a t
Answer #1

use Gauss' law,

E1 = Q/(A εo) = Q/(4πr² εo)

E2 = q/(4πr² εo)

E_tot = E1 + E2 = (Q + q)/(4πr² εo) = (8.75 + 3.03)E-6/(4π(0.199)² 8.85 E-12)

E_tot = 2676123.967 N/C